Shift report 20000721 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 21 2000 - 00:05:58 EDT

  • Next message: http account: "Shift report 20000721 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Jul 21 00:05:58 2000, Shift: 16:00-24:00
    Supervisor: YKL, with: TS
    Log book#3 pages 35 - 37
    
    Shift Summary:
    ------------------------------------------------------------
    Injection studies most of the time. There were stable
    beam at 65G, cogged, for 40 minutes, which yielded
    altogather a dozen collisions, see run 1861-2. 
    MTPC1, MTPC2, T1 and a number of Drift Chamber anodes
    tripped. TILE1 also tripped. Utilizing so many no-beam
    periods, cosmic ray calibration of MULTI tiles was
    carried out piecemeal.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 21 2000 - 00:09:32 EDT